# ARTalk: Speech-Driven 3D Head Animation via Autoregressive Model

### v4

@@ Abstract.
Speech-driven 3D facial animation aims to generate realistic lip movements and facial expressions for 3D head models from arbitrary audio clips. Although existing diffusion-based methods are capable of producing natural motions, their slow generation speed limits their application potential. In this paper, we introduce a novel autoregressive model that achieves real-time generation of highly synchronized lip movements and realistic head poses and eye blinks by learning a mapping from speech to a multi-scale motion codebook. Furthermore, our model can adapt to unseen speaking styles, enabling the creation of 3D talking avatars with unique personal styles beyond the identities seen during training. Extensive evaluations and user studies demonstrate that our method outperforms existing approaches in lip synchronization accuracy and perceived quality. Demos and codes are available at https://xg-chu.site/project_artalk/.
@@
copyright: nonejournal: TOG
@@ 2. Related Work
Figure 2. ARTalk involves two separated parts. (a) shows our temporal multi-scale VQ autoencoder. It encodes motion sequences into multi-scale token maps $[M_{k_{1}},M_{k_{2}},...,M_{K}]$ using a shared codebook and causal masking on temporal. (b) shows The ARTalk Causal Transformer, where training uses ground truth tokens with a block-wise causal attention mask, and inference autoregressively predicts motion tokens conditioned on speech features and last scale tokens and last time window motions.
@@ 2.3. Diffusion based 3D Facial Motion Generation
Diffusion models (Ho et al., 2020; Zhu et al., 2023; Alexanderson et al., 2023; Stan et al., 2023; Ma et al., 2024; Sun et al., 2024) have recently gained traction for generative tasks because of their strong modeling capacities. FaceDiffuser (Stan et al., 2023) conditions a diffusion model on audio features to predict displacements from a neutral template. More recently, FaceTalk (Aneja et al., 2024) employs a diffusion model on expression coefficients with cross-attention to wav2vec2 (Baevski et al., 2020) features, while DiffSpeaker (Ma et al., 2024) utilizes biased conditional attention for a diffusion-driven synthesis pipeline. Scantalk (Nocentini et al., 2024) uses a DiffusionNet (Sharp et al., 2022) structure to overcome the fixed topology limitation and perform diffusion on arbitrary meshes. Similarly, DiffPoseTalk (Sun et al., 2024) uses a diffusion-based transformer decoder to generate both expression and pose (blendshapes) conditioned on audio and learned style features. Going beyond motion, LetsTalk (Zhang et al., 2024) adopts a spatiotemporal diffusion approach to synthesize entire video frames. Although these methods generate high-fidelity and realistic facial motion, their iterative sampling steps can be computationally expensive and may limit real-time applicability.
In this paper, we leverage FLAME (Li et al., 2017) as our motion representation and propose a novel autoregressive framework for speech-driven motion generation. It not only outperform current diffusion models in generating facial motion, but it also achieves real-time speed and produce natural head poses. A comparison of our method with existing approaches is shown in Table 1. Our method enables arbitrary style inference, allowing stylized motion generation at inference time by providing an example motion clip. Additionally, our method generates both facial expressions and head poses while maintaining real-time performance.
@@ 3.1. Preliminaries
We define the motion vector $\mathbf{M}\in\mathbb{R}^{K\times D}$ as the concatenation of per-frame expression and pose parameters over $K$ frames, i.e., $\mathbf{M}=[\psi\;\;\theta]$. The corresponding vertex coordinates $\mathbf{V}\in\mathbb{R}^{K\times N\times 3}$ store the 3D positions of $N$ mesh vertices per frame, where $\mathbf{V}_{\mathrm{lips}}\subset\mathbf{V}$ denotes the subset corresponding to the lip region, which is particularly important for speech related expression modeling. Further details on 3DMM can be found in the supplementary material.
@@ 3.2. Temporal Multi-scale VQ Autoencoder
Predicting motion frames from speech is a discrete task involving highly dense temporal sequences and complex mappings. Inspired by the success of discrete representations in image and motion generation (Xing et al., 2023; van den Oord et al., 2017; Zhou et al., 2022) and the effectiveness of multi-scale encoding in (Tian et al., 2024), we propose a temporal multi-scale VQ autoencoder to efficiently model motion dynamics.
@@
Temporal causal reasoning. To ensure stability over extended sequences, we incorporate a causal reasoning mechanism that enforces cross-window consistency. Instead of generating frames independently, our approach considers two consecutive time windows, $T-1$ and $T$. During training and inference, causal masks are applied, ensuring that predictions for $T$ depend only on previously observed information from $T-1$, without leaking future details. This prevents temporal discontinuities and enhances the smoothness of long-term motion.
@@ 3.3. Speech-to-Motion Autoregressive Model
After training the VQ autoencoder (Section 3.2), we obtain discrete codes $\{\mathbf{z}_{T}^{(l)}\}$ at multiple resolutions $k_{l}$ for each time window $T$, where $l=1,\dots,L$ spans from coarsest $(k_{1})$ to finest $(k_{L})$. To ensure long-term temporal coherence, we model these codes with a Transformer that is autoregressive across scales $(1\!\to\!L)$ and across adjacent windows $(T-1\!\to\!T)$.
@@
Figure 3. Comparison of runtime efficiency and performance across different methods.
@@ 4. Experiments
Datasets. We use the TFHP dataset proposed by  Sun et al. (2024) to train our model, which consists of 1,052 video clips from 588 subjects, with a total duration of approximately 26.5 hours. All videos are tracked at 25 fps, resulting in approximately 2,385,000 action frames. For training and evaluation, we adopt the train/test split provided in the original paper (Sun et al., 2024). Furthermore, we evaluate the generalization performance of our model on the test split of the widely used VOCASET dataset (Cudeiro et al., 2019). The test split consists of 80 audiovisual sequence pairs from 2 subjects. To adapt VOCASET to our method, we track the FLAME parameters at 25 fps and conduct evaluation on our tracked data. It is worth noting that we don’t use the VOCASET training split to train or fine-tune our method.
Implementation Details. Our method is implemented on the PyTorch framework (Paszke et al., 2017). In the first stage, we train our VQ autoencoder to obtain a multi-scale motion codebook. The size of the motion codebook is 256 and the code dimension is 64. The motion codebook consists of 256 entries, each with a code dimension of 64. The time window size is 100 frames (4 seconds), and the multi-scale levels are [1, 5, 25, 50, 100]. During this stage, we used the AdamW optimizer with a learning rate of 1.0e-4, a total batch size of 64, and trained for 50,000 iterations. In the second stage, we train the multi-scale autoregressive model using the AdamW optimizer with the same learning rate of 1.0e-4 and a batch size of 64 for 50,000 iterations. During this stage, we employ a frozen pre-trained HuBERT (Hsu et al., 2021) backbone without fine-tuning.
@@ 4.1. Quantitative Results
Based on previous studies (Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Nocentini et al., 2024; Sun et al., 2024), we adopted two quantitative metrics, the lip vertex error (LVE) (Richard et al., 2021) and the upper face dynamic deviation (FDD) (Xing et al., 2023), to evaluate the generated facial motions. LVE calculates the maximum L2 error of all lip vertices for each frame, evaluating the largest deviation between predicted and ground truth lip positions. FDD calculates the standard deviation of the motion of each upper facial vertex over time between predictions and ground truth, evaluating the consistency of upper facial motion, which is closely related to speaking styles. Additionally, we use a similar metric as  Sun et al. (2024), mouth opening distance (MOD), to more accurately assess the stylistic similarity of mouth opening movements. MOD measures the average difference in mouth opening region between predictions and ground truth. Compared to LVE, MMD focuses more on the similarity of mouth opening styles and is less sensitive to temporal lip synchronization. For the partitioning of the lip area and upper face, we used the mask partitioning officially provided by FLAME (Li et al., 2017). The lip region contains 254 points, while the upper face, including eye region and forehead, contains 884 points.
@@
We present the quantitative comparison on TFHP dataset (Sun et al., 2024) in Table 2. The baseline methods FaceFormer (Fan et al., 2022), CodeTalker (Xing et al., 2023), SelfTalk (Peng et al., 2023a), and FaceDiffuser (Stan et al., 2023) are all mesh-based approaches and cannot directly generalize to arbitrary styles or new meshes. To ensure a fair comparison, we retrained these methods on meshes generated from the TFHP dataset. For MultiTalk (Sung-Bin et al., 2024), which is trained on FLAME meshes and supports language-based stylization, we directly used its English style for evaluation. ScanTalk claims to work on arbitrary meshes, so we use the officially provided pre-trained weights. For DiffPoseTalk, we directly used its pre-trained weights on the TFHP dataset (including head pose) for evaluation. The results in the Table 2 show that our method achieving significant improvements in lip synchronization accuracy (LVE) and style alignment (FFD and MOD), indicating that our method not only achieves precise lip synchronization but also effectively captures personalized speaking styles.
@@
To demonstrate the generalizability of our method, we evaluated it on the VOCASET dataset (Cudeiro et al., 2019). The results are presented in Table 3. Notably, for some baseline methods (Fan et al., 2022; Xing et al., 2023; Peng et al., 2023a; Stan et al., 2023; Nocentini et al., 2024), we used their pre-trained weights on VOCASET, whereas our method was not trained or fine-tuned on this dataset. Despite this, our method achieved highly competitive performance and outperformed most baseline methods specifically trained on this dataset. For MultiTalk (Sung-Bin et al., 2024) and DiffPoseTalk (Sun et al., 2024) which designed with style generalization capabilities, we followed the same strategy as with our method by directly testing it on the VOCASET dataset without any additional training or fine-tuning. The results indicate that our method surpasses them in terms of generalization ability, further verifying the robustness of our method in handling unseen styles and data.
@@ 4.2. Qualitative Results
Table 4. User study results.
Method Sync (%) Nat-Exp (%) Style (%) Nat-Pose (%)
vs. FaceFormer 75.0 89.3 88.1 -
vs. CodeTalker 84.5 86.9 92.9 -
vs. SelfTalk 85.7 89.3 91.7 -
vs. FaceDiffuser 88.1 90.4 86.9 -
vs. MultiTalk 78.6 76.2 78.6 -
vs. ScanTalk 88.1 90.5 90.5 -
vs. DiffPoseTalk 63.1 59.5 60.7 58.3
@@ 4.3. User Study
User studies are a reliable approach for evaluating 3D talking heads. To comprehensively compare our method with baseline methods (Fan et al., 2022; Xing et al., 2023; Peng et al., 2023a; Stan et al., 2023; Sung-Bin et al., 2024; Nocentini et al., 2024; Sun et al., 2024), we conducted a user study focusing on four key metrics: lip synchronization, facial expression naturalness, style consistency and head pose naturalness.
@@ 4.4. Ablation study
Key Hyperparameter Choices. We explored different number of motion frames within a single time window, which is critical for our method. Specifically, we tested window lengths of 8, 25, 50 frames. For a window length of 8, we used a multi-scale sequence of [1,2,4,8], while for 25 and 50, we adopted [1,5,10,15,25] and [1,5,10,25,50], respectively. The results, presented in Table 5, demonstrate that the choice of window size significantly impacts performance. However, the window size depends heavily on the downstream task characteristics, such as whether the input consists of streaming audio chunks. In general, longer temporal windows generally produce better results. In this paper, we chose a window length of 100 frames and a multi-scale sequence of [1, 5, 25, 50, 100] as it is consistent with previous work such as DiffPoseTalk (Sun et al., 2024) and achieves a good balance between quality and efficiency.
@@
Motion Clip Length 8 11.73 24.89 2.25
Motion Clip Length 25 10.20 19.00 1.97
Motion Clip Length 50 9.78 18.03 1.89
@@ 5. Discussion and Conclusion
Limitations and future work. While ARTalk demonstrates strong performance in lip synchronization and expressions, a current limitation is that head motions are primarily driven by speech prosody (rhythm and emphasis) rather than semantic context. Addressing the cultural and semantic nuances of head gestures, such as nodding or shaking to indicate affirmation, would require broader and more diverse datasets and fine-grained head pose control, which we leave as future work. We hope that our work lays a solid foundation for the further research of 3D talking head generation.
@@
[…]
Qualitative comparison.

### v5

@@ Abstract.
Speech-driven 3D facial animation aims to generate realistic lip movements and facial expressions for 3D head models from arbitrary audio clips. Although existing diffusion-based methods are capable of producing natural motions, their slow generation speed limits their application potential. In this paper, we introduce a novel autoregressive model that achieves real-time generation of highly synchronized lip movements and realistic head poses and eye blinks by learning a mapping from speech to a multi-scale motion codebook. Furthermore, our model can adapt to unseen speaking styles, enabling the creation of 3D talking avatars with unique personal styles beyond the identities seen during training. Extensive evaluations and user studies demonstrate that our method outperforms existing approaches in lip synchronization accuracy and perceived quality.
@@
submissionid: 1975journal: TOGjournalyear: 2025copyright: acmlicensedconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763955isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Computer graphics
@@ 2. Related Work
Figure 2. ARTalk involves two separated parts. (a) shows our temporal multi-scale VQ autoencoder. It encodes motion sequences into multi-scale token maps $M_{k_{1}},M_{k_{2}},...,M_{K}]$ using a shared codebook and causal masking on temporal. (b) shows The ARTalk Causal Transformer, where training uses ground truth tokens with a block-wise causal attention mask, and inference autoregressively predicts motion tokens conditioned on speech features and last scale tokens and last time window motions.
@@ 2.3. Diffusion based 3D Facial Motion Generation
Diffusion models (Ho et al., 2020; Zhu et al., 2023; Alexanderson et al., 2023; Stan et al., 2023; Aneja et al., 2024; Baevski et al., 2020; Ma et al., 2024; Sun et al., 2024; Zhang et al., 2024) have recently gained traction for generative tasks because of their strong modeling capacities. FaceDiffuser (Stan et al., 2023) conditions a diffusion model on audio features to predict displacements from a neutral template. Scantalk (Nocentini et al., 2024) uses a DiffusionNet (Sharp et al., 2022) structure to overcome the fixed topology limitation and perform diffusion on arbitrary meshes. Similarly, DiffPoseTalk (Sun et al., 2024) uses a diffusion-based transformer decoder to generate both expression and pose (blendshapes) conditioned on audio and learned style features. Although these methods generate high-fidelity and realistic facial motion, their iterative sampling steps can be computationally expensive and may limit real-time applicability.
In this paper, we leverage FLAME (Li et al., 2017) as our motion representation and propose a novel autoregressive framework for speech-driven motion generation. It not only outperform current diffusion models in generating facial motion, but also achieves real-time speed and produces natural head poses. A comparison of our method with existing approaches is shown in Table 1. Our method enables arbitrary style inference, allowing stylized motion generation at inference time by providing an example motion clip.
@@ 3.1. Preliminaries
We define the motion vector $\mathbf{M}\in\mathbb{R}^{K\times D}$ as the concatenation of per-frame expression and pose parameters over $K$ frames. The corresponding vertex coordinates $\mathbf{V}\in\mathbb{R}^{K\times N\times 3}$ store the 3D positions of $N$ mesh vertices per frame, where $\mathbf{V}_{\mathrm{lips}}\subset\mathbf{V}$ denotes the subset corresponding to the lip region, which is particularly important for speech related modeling. Further details on 3DMM can be found in the supplementary material.
@@ 3.2. Temporal Multi-scale VQ Autoencoder
Predicting motion frames from speech is a discrete task involving highly dense temporal sequences and complex mappings. Inspired by the success of discrete representations in image and motion generation (Xing et al., 2023; van den Oord et al., 2017; Zhou et al., 2022), the effectiveness of multiscale audio modeling for animation generation (Jung et al., 2024) and the effectiveness of multi-scale encoding in (Tian et al., 2024), we propose a temporal multi-scale VQ autoencoder to efficiently model motion dynamics.
@@
Temporal causal reasoning. To ensure stability over extended sequences, we incorporate a causal reasoning mechanism that enforces cross-window consistency. Our approach considers two consecutive time windows, T1T-1 and $T$. During training and inference, causal masks are applied, ensuring that predictions for $T$ depend only on previously observed information from T1T-1, without leaking future details. This prevents temporal discontinuities and enhances the smoothness of long-term motion.
@@ 3.3. Speech-to-Motion Autoregressive Model
After training the VQ autoencoder (Section 3.2), we obtain discrete codes $\{\mathbf{z}_{T}^{(l)}\}$ at multiple resolutions $k_{l}$ for each time window $T$, where $l=1,\dots,L$ spans from coarsest to finest. To ensure long-term temporal coherence, we model these codes with a Transformer that is autoregressive across scales $(1\!\to\!L)$ and across adjacent windows $(T-1\!\to\!T)$.
@@
Figure 3. Comparison of efficiency and performance across different methods.
@@ 4. Experiments
Datasets. We use the TFHP dataset proposed by  Sun et al. (2024) to train our model, which consists of 1,052 video clips from 588 subjects, with a total duration of approximately 26.5 hours. All videos are tracked at 25 fps, resulting in approximately 2,385,000 action frames. For training and evaluation, we adopt the train/test split provided in the original paper (Sun et al., 2024). Furthermore, we evaluate the generalization performance of our model on the test split of the widely used VOCASET dataset (Cudeiro et al., 2019). The test split consists of 80 audiovisual sequence pairs from 2 subjects. To adapt VOCASET to our method, we track the FLAME parameters at 25 fps and conduct evaluation on our tracked data. It is worth noting that we don’t use VOCASET to train or fine-tune our method.
Implementation Details. Our method is implemented on the PyTorch framework (Paszke et al., 2017). In the first stage, we train our VQ autoencoder to obtain a multi-scale motion codebook. The motion codebook consists of 256 entries, each with a code dimension of 64. The time window size is 100 frames (4 seconds), and the multi-scale levels are [1, 5, 25, 50, 100]. At this stage, we used the AdamW optimizer with a learning rate of 1.0e-4 and a total batch size of 64 for 50,000 iterations. In the second stage, we train the multi-scale autoregressive model using the AdamW optimizer with the same learning rate of 1.0e-4 and a batch size of 64 for 50,000 iterations. During this stage, we employ a frozen pre-trained HuBERT (Hsu et al., 2021) backbone.
@@ 4.1. Quantitative Results
Based on previous studies (Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Nocentini et al., 2024; Sun et al., 2024), we adopted two quantitative metrics, the lip vertex error (LVE) (Richard et al., 2021) and the upper face dynamic deviation (FDD) (Xing et al., 2023), to evaluate the generated facial motions. LVE calculates the maximum L2 error of all lip vertices for each frame, evaluating the largest deviation between predicted and ground truth lip positions. FDD calculates the standard deviation of the motion of each upper facial vertex over time between predictions and ground truth, evaluating the consistency of upper facial motion, which is closely related to speaking styles. Additionally, we use a similar metric as  Sun et al. (2024), mouth opening distance (MOD), to more accurately assess the stylistic similarity of mouth opening movements. MOD measures the average difference in mouth opening region between predictions and ground truth. Compared to LVE, MOD focuses more on the similarity of mouth opening styles and is less sensitive to temporal lip synchronization. For the partitioning of the lip area and upper face, we used the mask partitioning officially provided by FLAME (Li et al., 2017). The lip region contains 254 points, while the upper face, including eye region and forehead, contains 884 points.
@@
We present the quantitative comparison on TFHP dataset (Sun et al., 2024) in Table 2. The baseline methods FaceFormer (Fan et al., 2022), CodeTalker (Xing et al., 2023), SelfTalk (Peng et al., 2023a), and FaceDiffuser (Stan et al., 2023) are mesh-based approaches and cannot generalize to arbitrary styles or new meshes. To ensure a fair comparison, we retrained these methods on meshes generated from the TFHP dataset. For MultiTalk (Sung-Bin et al., 2024), which is trained on FLAME meshes and supports language-based stylization, we used its English style for evaluation. ScanTalk (Nocentini et al., 2024) claims to work on arbitrary meshes, so we use the officially provided pre-trained weights. For UniTalker (Fan et al., 2024), we compute the metrics using the meshes generated by the generalized pivot identities.
For DiffPoseTalk (Sun et al., 2024), we used pre-trained weights on the TFHP dataset (including head pose) for evaluation. The results in the Table 2 show that our method achieving significant improvements in lip synchronization accuracy (LVE) and style alignment (FFD and MOD), indicating that our method not only achieves precise lip synchronization but also effectively captures personalized speaking styles.
@@
To demonstrate the generalizability of our method, we evaluated it on the VOCASET dataset (Cudeiro et al., 2019). The results are presented in Table 3. Notably, for some baseline methods (Fan et al., 2022; Xing et al., 2023; Peng et al., 2023a; Stan et al., 2023; Nocentini et al., 2024), we used their pre-trained weights on VOCASET, whereas our method was not trained or fine-tuned on this dataset. Despite this, our method achieved highly competitive performance and outperformed most baseline methods specifically trained on this dataset. For MultiTalk (Sung-Bin et al., 2024) and DiffPoseTalk (Sun et al., 2024) which designed with style generalization capabilities, we directly testing it on the VOCASET dataset without fine-tuning. The results indicate that our method surpasses them in terms of generalization ability, further verifying the robustness of our method in handling unseen styles and data.
@@ 4.2. Qualitative Results
Table 4. User study results. The percentages represent the proportion of users who preferred the our method over the baseline in each category. The values in brackets represent the 95% confidence interval, indicating that we are 95% confident that the true value of the performance metric lies within this range. Among them, Sync represents lip synchronization, Style represents the consistency of style, N-Exp and N-Pose represent the facial expression naturalness and head pose naturalness respectively.
Method Sync (%) N-Exp (%) Style (%) N-Pose (%)
FaceFormer 75.0 [70.4, 79.6] 89.3 [86.3, 92.8] 88.1 [85.0, 91.8] -
CodeTalker 84.5 [80.7, 88.4] 86.9 [83.3, 90.5] 92.9 [90.5, 95.9] -
SelfTalk 85.7 [82.0, 89.5] 89.3 [86.3, 92.8] 91.7 [89.1, 94.9] -
FaceDiffuser 88.1 [85.0, 91.8] 90.4 [87.3, 93.6] 86.9 [83.3, 90.5] -
MultiTalk 78.6 [74.5, 83.2] 76.2 [72.0, 81.0] 78.6 [74.5, 83.2] -
ScanTalk 88.1 [85.0, 91.8] 90.5 [87.7, 93.9] 90.5 [87.7, 93.9] -
UniTalker 81.8 [77.7, 86.0] 83.6 [79.7, 87.6] 81.3 [77.4, 85.7] -
DiffPoseTalk 63.1 [58.2, 68.5] 59.5 [54.3, 64.8] 60.7 [55.5, 65.9] 58.3 [53.1, 63.6]
@@ 4.3. User Study
User studies are a reliable approach for evaluating 3D talking heads. To comprehensively compare our method with baseline methods, we conducted a user study focusing on four key metrics: lip synchronization, facial expression naturalness, style consistency and head pose naturalness.
@@ 4.4. Ablation study
Key Hyperparameter Choices. We explored different number of motion frames within a single time window, which is critical for our method. Specifically, we tested window lengths of 8, 25, 50 frames. For a window length of 8, we used a multi-scale sequence of [1,2,4,8], while for 25 and 50, we adopted [1,5,10,15,25] and [1,5,10,25,50], respectively. The results, presented in Table 5, demonstrate that the choice of window size significantly impacts performance. However, the window size depends heavily on the downstream task characteristics, such as whether the input consists of streaming audio chunks. In general, longer temporal windows generally produce better results, but longer windows mean that more audio needs to be collected to output motion. In this paper, we chose a window length of 100 frames and a multi-scale sequence of [1, 5, 25, 50, 100] as it is consistent with previous work such as DiffPoseTalk (Sun et al., 2024) and achieves a good balance between quality and efficiency. Further exploration of the number of multi-scale layers and audio chunk lengths is provided in the supplementary material.
@@
Clip Length 8 Layer 5 11.73 24.89 2.25
Clip Length 25 Layer 5 10.20 19.00 1.97
Clip Length 50 Layer 5 9.78 18.03 1.89
@@ 5. Discussion and Conclusion
Limitations and future work. While ARTalk demonstrates strong performance in lip synchronization and expressions, a key limitation is that head movements are driven by speech prosody, not semantic context. This prevents the model from generating culturally-specific or meaningful gestures like nodding or shaking the head to indicate affirmation. Additionally, our use of a single style embedding can be challenging for speech with dramatic emotional shifts, which is also a common problem with existing methods. For example, when a speaker suddenly shifts from calm to excited, a single embedding may struggle to accurately capture this transition.
We hope that our work lays a solid foundation for the further research of 3D talking head generation. Future research will focus on addressing the above challenges, especially exploring how to effectively incorporate semantic information into head motion generation and developing more flexible style control methods.
@@
Acknowledgements.
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. In addition, this work was also partially supported by JST SPRING, Grant Number JPMJSP2108.